LASER SPECTROSCOPY OF THE Ã^{2}Π − X̃^{2}Σ^{+} TRANSITION OF BaOH

J.-G. WANG, J. D. TANDY, and \underline{P. F. BERNATH}, Department of Chemistry, University of York, Heslington, York, YO10 5DD, UK.

The Ã^{2}Π−X̃^{2}Σ^{+} transition of BaOH has been rotationally analyzed using the technique of high resolution laser excitation spectroscopy. The BaOH molecules were synthesized in a Broida-type oven and a single mode Ti:Sapphire laser was used to record the spectra. Two vibronic bands (001 − 001 and 001 − 000) in addition to the 000 − 000 band have also been observed and assigned. Rotational and fine structure parameters were determined through a combined least-squares fit with the millimeter-wave pure rotational data of the ground state.